write the equation of the line that satisfies the given conditions. Contains the point (-1,4) and is perpendicular to the line x-3y=3.
Answer by rothauserc(4718)

1) x -3y = 3
:
rewrite equation 1 in slope - intercept form
:
y = x/3 - 1
:
the negative reciprocal of the slope, is the slope of a line perpendicular to the line in equation 1
:
slope of perpendicular line is -3
:
2) y = -3x +b
:
now use the point (-1,4) by substituting for x and y in equation 2
:
4 = -3(-1) +b
:
b = 1
:
********************************************************************
equation of the line with the given conditions is
:
y = -3x +1 (slope - intercept form)
:
3x +y = 1 (standard form)
********************************************************************
: